OpenStudy (jrwarbington18):

Please Help! Simplify the expression below if x > 0 and y > 0. x√y/y√x a. √y/xy b. √y/y c. √xy/y d. x√xy/y

OpenStudy (eliesaab):

Is this your question Simplify \[ \frac{ x \sqrt y}{y \sqrt x} \]

OpenStudy (eliesaab):

How do you simplify \[ \frac x{\sqrt x} \]

OpenStudy (eliesaab):

\[ \frac x{\sqrt x}=\frac{ x^1}{x^{1/2}}=x^{1-1/2}=? \]

OpenStudy (eliesaab):

Do the same for \[ \frac{\sqrt y}y \]

OpenStudy (eliesaab):

Then combine and find the answer

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)